# When Simultaneous Localization and Mapping Meets Wireless Communications: A Survey

### v1

@@ Abstract
The availability of commercial wireless communication and sensing equipment combined with the advancements in intelligent autonomous systems paves the way towards robust joint communications and simultaneous localization and mapping (SLAM). Autonomous vehicles are one of the several applications that directly benefits from millimeter-wave (mmWave) access, enabling improved communications and accurate localization in the interest of control and navigation. This paper surveys the state-of-the-art in the nexus of SLAM and Wireless Communications, attributing the bidirectional impact of each with a focus on visual SLAM (V-SLAM) integration. We provide an overview of key concepts related to wireless signal propagation, geometric channel modeling, and radio frequency (RF)-based localization and sensing. In addition to this, we show image processing techniques that can detect landmarks, proactively predicting optimal geometric paths for wireless channels. Several dimensions are considered, including the prerequisites, techniques, background, and future directions and challenges of the intersection between SLAM and wireless communications. We analyze mathematical approaches such as Bayesian filters, probabilistic state estimation, motion models, vector fields, and spatial methods for signal processing, as well as key technological aspects. We expose techniques and items towards enabling a highly effective retrieval of the autonomous robot/vehicle kinematic state. Among other interesting findings, we observe that monocular V-SLAM would greatly benefit from RF/mmWave relevant information, as the latter can serve as a proxy for the scale ambiguity resolution. Conversely, we find that wireless communications in the context of 5G and beyond can potentially benefit from the visual odometry and 6-D pose estimation methods that are central in SLAM. Moreover, we examine other perception sources besides the camera for SLAM such as Radio SLAM, wireless fidelity (WiFi) SLAM, and light detection and ranging (LiDAR) SLAM, and describe the twofold relation with wireless communications. Finally, integrated solutions performing joint communications and localization are still in their infancy: theoretical and practical advancements are required to add higher-level localization and semantic perception capabilities to RF/mmWave and multi-antenna technologies.
@@ I Introduction
On the other hand, wireless communications have undergone several impactful advancements, such as high-frequency and ultra-wide bandwidth capabilities, reduced latency, and increased coverage and reliability [108]. The advent of networks endowed with these capabilities has provided the physical foundation for this transformation, enabling wireless signals to convey both data and perceptual information. Among these, mmWave technologies, operating between 28 and 300 GHz, stand out for their ability to support dense connectivity and precise spatial resolution [128]. As mmWave technology continues to evolve, it is deemed a fundamental enabler of sixth-generation (6G) mobile communication networks, which are expected to be commercially deployed around 2030. Driven by the need to support emerging applications such as connected autonomous vehicles, connected robots, and mixed reality, with all these applications exceeding the capabilities of current fifth-generation (5G) networks, researchers have outlined the 6G vision. 6G will aim at global coverage, considerably wide spectrum and strong security [141] as well as a single intelligent infrastructure where communication, sensing, learning, and control converge. Recently, the International Telecommunication Union (ITU) introduced a draft recommendation defining the framework and overarching objectives for International Mobile Telecommunications-2030 (IMT-2030) and beyond, highlighting six key usage scenarios: immersive communication, massive communication, hyper-reliable and low-latency communication (HRLLC), ubiquitous connectivity, and the integration of advanced artificial intelligence (AI) algorithms with communications [148].
@@
Next-generation wireless communication technology serves communication purposes while also providing high-resolution user equipment (UE)-based localization and device-free sensing. These capabilities are inherent to the physics of mmWave propagation. To begin with, the shorter wavelength of mmWaves, compared to sub-6G signals, aids in low error-induced position estimation. Additionally, mmWave sensing results in higher spatial scanning resolution due to propagation characteristics such as dominance of the line-of-sight (LOS) propagation component over the scattered ones that follow from propagation in areas with many reflecting surfaces [108]. Nevertheless, mmWave propagation is capable of long-range sensing, the latter at the cost of higher path loss and attenuation compared to lower frequency radio-waves. This can be tackled by: a) employing large or massive antenna arrays and b) devising specific protocols for initial access as well as beam training in the context of IEEE 802.11ay standard, in order for the electromagnetic field to be focused on a desired region in the cartesian space [153].
@@
Device-free localization (DFL), also known as radio sensing or sensor-less sensing, determines object locations by analyzing wireless link measurements. The key wireless properties used in this technique are received signal strength (RSS) and channel state information (CSI). DFL methods fall into two main categories based on how they utilize the aforementioned measurements [110]. The first one, link quality measurement, assesses changes in signal strength caused by objects obstructing the wireless link. A probabilistic/stochastic model for the three-dimensional position state prior distribution can be combined with RSS measurement updates to estimate state posterior distribution, adopting a simplified odometry assumption such as a constant-velocity model. The second category, titled reflection or scattering of the link, relies on signal reflections and scattering from objects to estimate their location, making it analogous to a radar-based approach [110]. A typical scenario would include a number of transmitter-receiver pairs operating as a network of sensor radars, whilst in addition to the target objects, background objects are considered. Multi-object detection and localization can be established by placing nodes with ultra-wideband (UWB) sensing capabilities in the vicinity of the region of interest for target detection. The computation can be based on the time of arrival (ToA) of the back-scattered signals.
@@
Although numerous recent surveys have investigated various aspects of wireless localization and sensing, few have examined the bidirectional nexus between SLAM or its core computer vision related elements and wireless communications. Among the most closely related works, Zeng et al. [141] provided a tutorial on environment-aware communications through the concept of channel knowledge maps (CKM), highlighting fingerprint-based localization and integrated sensing–communication frameworks. However, it remains limited to the cellular-centric extraction of positioning and spatial information without extending into autonomous SLAM or computer vision-aided wireless communication aspects. Italiano et al. [46] presented a comprehensive tutorial on 5G capabilities, detailing THz-band imaging and high-frequency SLAM for highly accurate positioning and NLOS scenarios leveraging multipath reflections. Moreover, they included in their work reconfigurable intelligent surfaces (RIS)-aided holographic localization and probabilistic SLAM-inspired data association, yet their treatment remains confined to UE localization accuracy and does not unify mathematical SLAM foundations or vision-driven autonomy. Similarly, Shastri et al. [108] and Cheraghinia et al. [21] covered device-based and device-free mmWave and UWB radar localization, respectively, offering valuable insights into channel-level processing, angle-of-arrival estimation, and tracking. However, their work did not involve an integral perspective incorporating such modalities into broader SLAM frameworks involving multi-sensor fusion, control, and learning-based mapping.
@@
Beyond these highly overlapping works, moderately related surveys such as Primeau et al. [92], Moura et al. [80], and Kong et al. [52] examined computational intelligence-based wireless sensor and actuator networks, opportunistic localization for sensor networks, and mmWave Radar based sensing with a focus on machine learning techniques, respectively. However, each study addressed these mechanisms in an uncombined computer vision and wireless communications SLAM setting. Other reviews such as Liu et al. [71] focused on integrated sensing and communication (ISAC) limits, highlight the rising interest in environment-aware communication, yet none have connected these to the algorithmic or autonomy-enabling layers of SLAM. In contrast, the present survey synthesizes these research threads into a unified taxonomy encompassing (i) the foundations of autonomous SLAM, and (ii) mathematical methods and multimodal SLAM techniques, while highlighting the transformative role of modern wireless communication technologies in the era of 6G, ranging from WiFi fingerprinting and radio map construction to deep neural network (DNN)–based wireless SLAM and RIS-aided MetaSLAM. This integrative treatment establishes a cross-disciplinary reference point for understanding how wireless signals are increasingly redefining localization, mapping, and situational awareness across next-generation robotic and communication systems. Table I compares this survey with closely related works in terms of their treatment of SLAM as a core framework, joint consideration of vision-wireless communications and the integration of perception, control, and communications within a combined approach.
Specifically, this survey provides a structured and in-depth investigation of the emerging intersection of SLAM and wireless communications, offering a novel taxonomy that captures both foundational and new developments. Anchored around three core thematic axes: the foundations of SLAM in networked autonomous systems, SLAM methods and techniques, and the advancements of wireless communications in SLAM. In the aforementioned areas of focus, a comprehensive synthesis of the state-of-the-art is provided. Our study contextualizes the pillars for networked autonomous systems’ SLAM, highlighting the critical role of the computer vision algorithms for extracting features that can serve as the basis for reconstruction of the environment as well as of the control laws in autonomous and connected systems. Based on these pillars, the paradigm of passive SLAM is highlighted within the larger picture of autonomous systems, where SLAM has the responsibility of being a state observer using diverse radio and visual features. Optimization of the control inputs in order for a networked autonomous system to reduce its ego-motion or landmark mapping uncertainty is considered afterwards, contextualizing the domain of active estimation or active SLAM. The latter paradigm can be central to a pipeline for achieving communication goals, i.e., by conversion of the wireless communications’ objective such as path-loss reduction/RSS improvement through optimal beam steering to a perceptual goal. Thus, a goal such as improving the accuracy of UE position and orientation through the execution of specific motion commands can be directly translated into a beam prediction objective.
On the methodological front, the survey focuses on mathematical formulations, RF and computer vision–based localization paradigms, LiDAR and inertial SLAM, and the algorithmic underpinnings that define modern SLAM systems. Furthermore, the survey foregrounds the integration of wireless technologies into SLAM pipelines, covering developments in device-based sensing, WiFi fingerprinting-based localization, radio map construction via wireless sensing, and emerging DNN-driven wireless SLAM architectures. By organizing these contributions in a unified framework, the survey demonstrates how wireless signals augment the vehicle state vector and redefine localization and mapping strategies in an advantageous manner. As the wireless communications-computer vision intersection lives in the core of this study, effort is put into showcasing the importance of computer vision pipelines for predicting the environment state. This piece of information can be particularly useful for predicting future link quality.
The organization of the paper is as follows. Section II is dedicated to the historical background. Section III presents the foundations of SLAM in networked autonomous systems. SLAM methods, modalities and techniques are presented in Section IV. Section V discusses the challenges arising from the SLAM–wireless communications nexus and outlines future research directions. Finally, Section VI concludes the paper.
@@
TABLE I: Comparison of Representative Surveys Related to SLAM and Wireless Communications
@@
Survey Primary Focus Coverage of SLAM as Core Framework Vision-Wireless Communications Joint Consideration System-Level Perception-Control-Communications Approach
Zeng et al. [141] Environment-aware communications, CKM Limited (cellular-centric localization) No No
Italiano et al. [46] 5G positioning and sensing Partial (focuses on SLAM-inspired UE positioning) No No
Shastri et al. [108] mmWave device-based localization Limited (Radar-enabled localization and sensing, indoor AP-based localization) No No
Cheraghinia et al. [21] UWB radar sensing and tracking Limited (UWB radar localization, AoA estimation, tracking) No No
Primeau et al. [92] Computational intelligence in Wireless Sensor Actuator Networks Limited (localization presented as an input to actuators) No Partial (interconnections between sensing and actuation approached from an algorithmic perspective)
@@
Moura et al. [80] Opportunistic localization, Multi-Access Edge Computing No No No
@@
Kong et al. [52] Radar-based sensing Partial (mmWave sensing / localization specific) No No
This Survey SLAM–Wireless Communications Nexus Yes Yes (Vision + RF/mmWave) Yes (Passive & Active SLAM, Control-Aware)

### v2

@@ Abstract
The availability of commercial wireless communication and sensing equipment combined with the advancements in intelligent autonomous systems contribute to robust joint communications and simultaneous localization and mapping (SLAM). Autonomous vehicles are one of the several applications that directly benefit from millimeter-wave (mmWave) access, facilitating improved communications and accurate localization in the interest of control and navigation. This paper surveys the state-of-the-art in the intersection of SLAM and Wireless Communications, attributing the bidirectional impact of each with a focus on visual SLAM (V-SLAM) integration. We provide an overview of key concepts related to geometric channel modeling, and radio frequency (RF)-based localization and sensing. Several research directions are considered, including the interpretation of SLAM as a unified perception-estimation framework driven by wireless communications-based positioning and visual information, the integration of communication-aware objectives in active perception, and multi-modal odometry through the fusion of wireless measurements and complementary sensing modalities. We analyze estimation and control approaches such as Bayesian filters, feature-based pose estimation, perception-aware motion control, spatial methods for signal processing such as vector fields, and key technological aspects. Among other interesting findings, we observe that monocular V-SLAM might greatly benefit from RF/mmWave signals, as the latter can serve as a proxy for scale ambiguity resolution. Conversely, we find that wireless communications in the context of 5G and beyond can potentially benefit from visual odometry and landmark detection methods in the context of optimal geometric path prediction for wireless channels. Moreover, we examine other perception sources besides monocular vision for SLAM such as Radio SLAM, wireless fidelity (WiFi) SLAM, and light detection and ranging (LiDAR) SLAM, and describe a twofold relation with wireless communications. Finally, integrated solutions performing joint communications and localization appear to be in their infancy: theoretical and practical advancements could add higher-level localization and semantic perception capabilities to RF/mmWave and multi-antenna technologies.
@@ I Introduction
On the other hand, wireless communications have undergone several impactful advancements, such as high-frequency and ultra-wide bandwidth capabilities, reduced latency, and increased coverage and reliability [110]. The advent of networks endowed with these capabilities has provided the physical foundation for this transformation, enabling wireless signals to convey both data and perceptual information. Among these, mmWave technologies, operating between 28 and 300 GHz, have the potential to support dense connectivity and precise spatial resolution [130]. As mmWave technology continues to evolve, it is deemed a contributor of sixth-generation (6G) mobile communication networks, which are expected to be commercially deployed around 2030. Driven by the need to support emerging applications such as connected autonomous vehicles, connected robots, and mixed reality, with all these applications exceeding the capabilities of current fifth-generation (5G) networks, researchers have outlined the 6G vision. 6G will aim at global coverage, considerably wide spectrum and strong security [143] as well as a single intelligent infrastructure where communication, sensing, learning, and control converge. Recently, the International Telecommunication Union (ITU) introduced a draft recommendation defining the framework and overarching objectives for International Mobile Telecommunications-2030 (IMT-2030) and beyond, highlighting six key usage scenarios: immersive communication, massive communication, hyper-reliable and low-latency communication (HRLLC), ubiquitous connectivity, and the integration of advanced artificial intelligence (AI) algorithms with communications [151].
@@
Next-generation wireless communication technology serves communication purposes while also providing high-resolution user equipment (UE)-based localization and device-free sensing. These capabilities are inherent to the physics of mmWave propagation. To begin with, the shorter wavelength of mmWaves, compared to sub-6G signals, aids in low error-induced position estimation. Additionally, mmWave sensing can result in higher spatial scanning resolution due to propagation characteristics such as dominance of the line-of-sight (LOS) propagation component over the scattered ones that follow from propagation in areas with many reflecting surfaces [110]. Nevertheless, mmWave propagation has the potential to achieve long-range sensing, the latter at the cost of higher path loss and attenuation compared to lower frequency radio-waves. This can be tackled by: a) employing large or massive antenna arrays and b) devising specific protocols for initial access as well as beam training in the context of IEEE 802.11ay standard, in order for the electromagnetic field to be focused on a desired region in the cartesian space [156].
@@
Device-free localization (DFL), also known as radio sensing or sensor-less sensing, can retrieve object locations by analyzing wireless link measurements. The key wireless properties used in this technique are received signal strength (RSS) and channel state information (CSI). DFL methods fall into two main categories based on how they utilize the aforementioned measurements [112]. The first one, link quality measurement, assesses changes in signal strength caused by objects obstructing the wireless link. A probabilistic/stochastic model for the three-dimensional position state prior distribution can be combined with RSS measurement updates to estimate state posterior distribution, adopting a simplified odometry assumption such as a constant-velocity model. The second category, titled reflection or scattering of the link, relies on signal reflections and scattering from objects to estimate their location, making it analogous to a radar-based approach [112]. A typical scenario would include a number of transmitter-receiver pairs operating as a network of sensor radars, whilst in addition to the target objects, background objects are considered. Multi-object detection and localization can be established by placing nodes with ultra-wideband (UWB) sensing capabilities in the vicinity of the region of interest for target detection. The computation can be based on the time of arrival (ToA) of the back-scattered signals.
@@
Although several recent surveys have investigated various aspects of wireless localization and sensing, few have examined the bidirectional nexus between SLAM or its core computer vision related elements and wireless communications. Among the most closely related works, Katare et al.[50] reviewed the foundational elements of automated driving, assigning each to the related sense–think–act functionality as well as autonomous-driving perception, SLAM, and vehicle-to-everything (V2X) communications. However, their main focus was energy-efficient approximate edge AI, not the bidirectional SLAM–wireless estimation and control interplay. Zeng et al. [143] provided a tutorial on environment-aware communications through the concept of channel knowledge maps (CKM), highlighting fingerprint-based localization and integrated sensing–communication frameworks. However, it remains limited to the cellular-centric extraction of positioning and spatial information without extending into autonomous SLAM or computer vision-aided wireless communication aspects. Italiano et al. [48] presented a comprehensive tutorial on 5G capabilities, detailing THz-band imaging and high-frequency SLAM for highly accurate positioning and NLOS scenarios leveraging multipath reflections. Moreover, they included in their work reconfigurable intelligent surfaces (RIS)-aided holographic localization and probabilistic SLAM-inspired data association, yet their treatment remains confined to UE localization accuracy and does not unify mathematical SLAM foundations or vision-driven autonomy. Similarly, Shastri et al. [110] and Cheraghinia et al. [22] covered device-based and device-free mmWave and UWB radar localization, respectively, offering valuable insights into channel-level processing, angle-of-arrival estimation, and tracking. However, their work did not involve an integral perspective incorporating such modalities into broader SLAM frameworks involving multi-sensor fusion, control, and learning-based mapping.
@@
Beyond these highly overlapping works, moderately related surveys such as Primeau et al. [94], Moura et al. [82], and Kong et al. [54] examined computational intelligence-based wireless sensor and actuator networks, opportunistic localization for sensor networks, and mmWave Radar based sensing with a focus on machine learning (ML) techniques, respectively. However, each study addressed these mechanisms in an uncombined computer vision and wireless communications SLAM setting. Other reviews such as Liu et al. [73] focused on integrated sensing and communication (ISAC) limits, highlight the rising interest in environment-aware communication, yet none have connected these to the algorithmic or autonomy-supporting layers of SLAM. In comparison, the present survey synthesizes these research threads into a unified framework comprising (i) the foundations of autonomous SLAM, and (ii) mathematical methods and multimodal SLAM techniques, while highlighting the transformative role of modern wireless communication technologies in the era of 6G, ranging from WiFi fingerprinting and radio map construction to deep neural network (DNN)–based wireless SLAM and RIS-aided MetaSLAM. This integrative treatment can provide a cross-disciplinary reference point for understanding how wireless signals are increasingly redefining localization, mapping, and situational awareness across next-generation robotic and communication systems.
The literature search was conducted using a three-step approach to ensure broad coverage of relevant studies. First, a search was performed in the Scopus database using predefined keywords “(SLAM OR Simultaneous Localization and Mapping) AND (5G OR 6G OR WiFi OR Wireless Communications*). The search was limited to articles published up to September 2025. Second, a targeted manual search was conducted of all surveys published over the period January 2018 – September 2025 in IEEE Communications Surveys & Tutorials, to capture recent surveys in the field. This step was undertaken to ensure inclusion of influential and field-specific syntheses that may not have been fully captured through the Scopus database indexing. Finally, reference lists of the identified articles/surveys were screened to identify additional relevant studies and reports not retrieved through the latter search.
Table I compares this survey with closely related works in terms of the following dimensions: (i) SLAM as a core framework, (ii) joint consideration of vision - wireless communications, (iii) inclusion of mathematical models for estimation, and (iv) passive and (v) active SLAM within a combined perception - communications - control approach.
@@
TABLE I: Comparison of Representative Surveys and Tutorials Related to SLAM and Wireless Communications
@@
Survey Primary Focus SLAM as a Core Framework Joint Vision-Wireless Mathematical Models Passive SLAM Active SLAM
Zhou et al. [156] IEEE 802.11ay mmWave WLAN design, beam training No Low No No No
del Peral-Rosado et al. [26] Cellular mobile radio localization from 1G to 5G No Low No Moderate No
Primeau et al. [94] Computational intelligence in wireless sensor and actuator networks No Low No Yes High
@@
Moura and Hutchison [82] Game theory for multi-access edge computing, opportunistic localization No Low Moderate No No
Shit et al. [112] Device-free localization taxonomy for smart environments No Low High High No
Liu et al. [73] Fundamental limits of integrated sensing and communication No Low Yes Low No
Shastri et al. [110] mmWave device-based localization and device-free sensing No Low High High No
Zeng et al. [143] Environment-aware communications via CKM for 6G, environment-aware localization and sensing with a focus on mathematical models for estimation No Low Yes No No
Katare et al. [50] Energy-aware approximate edge AI, sense–think–act scheme in autonomous systems, perception, object detection, RGB / LiDAR visual SLAM, V2X communications Moderate High Low Yes Low
Italiano et al. [48] 5G positioning, mathematical models attributing different wireless measurement signals (e.g. ToA, AoA), UE positioning, mapping in harsh NLOS Moderate No Yes Low No
@@
Kong et al. [54] mmWave radar-based sensing applications and signal processing, focus on localization. Moderate No Moderate Low No
Cheraghinia et al. [22] UWB radar applications, signal processing, datasets, and trends Moderate Moderate High Yes No
This Survey SLAM - wireless communications integration in autonomous systems Yes Yes Yes Yes Yes
The comparison was based on a scoring methodology presented in Table II. Each dimension is assessed using four defined binary checks, which are assigned a value of 1 only when clear supporting evidence is identified in the corresponding paper, and 0 otherwise. The total score per dimension was computed as the sum of its four checks, resulting in a state in $\{0/4,1/4,2/4,3/4,4/4\}$.
These states are mapped to qualitative levels as follows: No ($0/4$), Low ($1/4$), Moderate ($2/4$), High ($3/4$), and Yes ($4/4$), reflecting increasing degrees of alignment with the defining characteristics of each dimension.
TABLE II: Dimensions’ Scoring Criteria and Assessment Logic for the Comparison Framework
Dimension Criterion Check Binary Condition for Assignment Contribution
SLAM as a Core Framework SLAM central in title / abstract / introduction 1 Explicit mention of SLAM as a primary focus in title, abstract, or introduction 0–4
Mapping jointly treated with localization 2 Joint discussion of mapping and localization processes within the survey
SLAM components / pipelines discussed 3 Significant treatment of SLAM pipelines, modules, or algorithmic components
SLAM as the organizing lens 4 Survey structure is centered around SLAM concepts or taxonomy
Joint Vision - Wireless Vision modalities discussed 1 Inclusion of RGB, stereo, RGB-D or LiDAR vision-based sensing modalities 0–4
RF / mmWave communications discussed 2 Discussion of 5G / 6G, RF / mmWave (cellular or WiFi) communication systems
Joint analytical treatment 3 Vision and wireless systems are treated within a unified analytical / modeling framework
Explicit integration 4 Direct discussion of interaction or fusion between vision and wireless modalities
Mathematical Models Measurement / signal / state-space models 1 Presence of explicit equations, signal models, or state-space formulations 0–4
Estimation/probabilistic methods 2 Use of stochastic, Bayesian / probabilistic estimation techniques
Optimization/bounds 3 Central use of optimization problems, bounds, or objective functions
Analytical depth 4 Predominantly analytical treatment
Passive SLAM SLAM in perception - communications - control context 1 Localization / mapping embedded in perception – communication – control systems 0–4
Link to control/autonomy 2 State estimation explicitly connected to control or decision-making
System-level coupling 3 Integration across system layers rather than isolated modules
Passive formulation 4 No explicit action optimization or active information-seeking mechanism required
Active SLAM Control / action discussed in perception – communication – control context 1 Explicit treatment of control inputs or action selection mechanisms 0–4
Action improves estimation / communication 2 Actions linked to uncertainty reduction, mapping, localization, or communication performance
Active perception loop 3 Closed-loop perception or information-seeking behavior is present
Action-estimation coupling 4 Control integrated within estimation or sensing-aided communication loops
Scoring Rule: Each dimension is assessed using four binary checks. A check is assigned value 1 when supporting evidence is identified in the paper, otherwise, it is assigned conservatively 0. The total score per axis is computed as the sum of its four checks, yielding a state in $\{0/4,1/4,2/4,3/4,4/4\}$, which is mapped to the qualitative labels No, Low, Moderate, High, and Yes, respectively.
Specifically, this survey provides a structured and in-depth investigation of the emerging intersection of SLAM and wireless communications, organizing both foundational and new developments. Centered on three core thematic axes: the foundations of SLAM in networked autonomous systems, SLAM methods and techniques, and the advancements of wireless communications in SLAM. In the aforementioned areas of focus, an overview of the state-of-the-art is provided. Our study describes the pillars for networked autonomous systems’ SLAM, highlighting the critical role of the computer vision algorithms for extracting features that can serve as the basis for reconstruction of the environment as well as of the control laws in autonomous and connected systems. Based on these pillars, the paradigm of passive SLAM is highlighted within the larger picture of autonomous systems, where SLAM has the responsibility of being a state observer using diverse radio and visual features. Optimization of the control inputs in order for a networked autonomous system to reduce its ego-motion or landmark mapping uncertainty is considered afterwards, highlighting the domain of active estimation or active SLAM. The latter paradigm could be central to a pipeline for achieving communication goals, i.e., by conversion of the wireless communications’ objective such as path-loss reduction/RSS improvement through optimal beam steering to a perceptual goal. Thus, a goal such as improving the accuracy of UE position and orientation through the execution of specific motion commands could be directly translated into a beam prediction objective.
On the methodological perspectives, the survey focuses on mathematical formulations, RF and computer vision–based localization paradigms, LiDAR and inertial SLAM, and algorithmic elements in modern SLAM systems. Furthermore, the survey examines the integration of wireless technologies into SLAM pipelines, addressing developments in device-based sensing, WiFi fingerprinting-based localization, radio map construction via wireless sensing, and emerging DNN-driven wireless SLAM architectures. By organizing these contributions in a unified framework, the survey demonstrates how wireless signals augment the vehicle state vector and redefine localization and mapping strategies in an advantageous manner. As the wireless communications-computer vision intersection lives in the core of this study, effort is put into showcasing the importance of computer vision pipelines for predicting the environment state. This piece of information can be particularly useful for predicting future link quality.
The organization of the paper is as follows. Section II provides the historical background. Section III presents the foundations of SLAM in networked autonomous systems. SLAM methods and modality-based techniques are presented in Section IV. Section V discusses wireless- and 6G-enabled SLAM. Section VI discusses challenges and future directions. Finally, Section VII concludes the paper.